OpenStudy (anonymous):

The volume of a cone that fits exactly inside a cylinder is 20 cubic feet. What is the volume of the cylinder?

OpenStudy (anonymous):

@whpalmer4

jimthompson5910 (jim_thompson5910):

If a cone fits exactly inside a cylinder (both have the same radius and height), then the rule is this Volume of Cylinder = 3*(Volume of Cone)

OpenStudy (anonymous):

x=3*20?

jimthompson5910 (jim_thompson5910):

Basically this is saying that you can fit 3 of these cones in the cylinder.

jimthompson5910 (jim_thompson5910):

exactly, so 60 cubic feet

OpenStudy (anonymous):

can you help me with another one?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Using the following equation, find the center and radius: x2 − 4x + y2 + 8y = −4

OpenStudy (anonymous):

i know that you have to isolate them

jimthompson5910 (jim_thompson5910):

no you just need to get it into the form (x-h)^2 + (y-k)^2 = r^2

jimthompson5910 (jim_thompson5910):

you do that by completing the square for x and y

OpenStudy (anonymous):

i dont get it

jimthompson5910 (jim_thompson5910):

what is the x coefficient?

OpenStudy (anonymous):

2 and -4

jimthompson5910 (jim_thompson5910):

2 is the exponent

jimthompson5910 (jim_thompson5910):

-4 is the coefficient for the x term

jimthompson5910 (jim_thompson5910):

cut that in half to get what?

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

square that to get ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

now add that to both sides

jimthompson5910 (jim_thompson5910):

Doing that gives you x^2 - 4x + y^2 + 8y = -4 x^2 - 4x + y^2 + 8y + 4 = -4+4 x^2 - 4x + y^2 + 8y + 4 = 0

jimthompson5910 (jim_thompson5910):

now if you rearrange terms, you would get this x^2 - 4x + y^2 + 8y + 4 = 0 (x^2 - 4x + 4) + y^2 + 8y = 0 what do you get when you factor x^2 - 4x + 4 ?

OpenStudy (anonymous):

(x−2)^2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

we now have (x-2)^2 + y^2 + 8y = 0

jimthompson5910 (jim_thompson5910):

so we're halfway done

jimthompson5910 (jim_thompson5910):

do the same thing for the y terms find the y coefficient, take half of that, then square that to get ???

OpenStudy (anonymous):

coefficients : 1,8

jimthompson5910 (jim_thompson5910):

not the y^2 term, just the y term

jimthompson5910 (jim_thompson5910):

sorry I should have clarified

OpenStudy (anonymous):

but i dont know how to take half

jimthompson5910 (jim_thompson5910):

what is the y coefficient? of the y term, not the y^2 term

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

cut that in half, then square that result

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

add that to both sides

jimthompson5910 (jim_thompson5910):

you'll have y^2 + 8y + 16 which factors to what?

OpenStudy (anonymous):

(y+4)^2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

we now have (x-2)^2 + (y+4)^2 = 16

jimthompson5910 (jim_thompson5910):

what is the center and radius?

OpenStudy (anonymous):

Radius: 16 Center: (-2,4)

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

keep in mind that the center of (x-h)^2 + (y-k)^2 = r^2 is (h,k) the radius of (x-h)^2 + (y-k)^2 = r^2 is r

OpenStudy (anonymous):

Center: (−2, −4) Radius: 16.

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

@jim_thompson5910 you there?

jimthompson5910 (jim_thompson5910):

(x-2)^2 + (y+4)^2 = 16 turns into (x-2)^2 + (y - (-4))^2 = 4^2

jimthompson5910 (jim_thompson5910):

notice how I'm replacing y + 4 with y - (-4). Think of this as two negatives canceling to make a positive

jimthompson5910 (jim_thompson5910):

and I'm rewriting 16 as 4^2

jimthompson5910 (jim_thompson5910):

now list off: h, k, r

jimthompson5910 (jim_thompson5910):

use (x-h)^2 + (y-k)^2 = r^2 as a template

OpenStudy (anonymous):

so what i put is not right?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

the radius is 4?

jimthompson5910 (jim_thompson5910):